From: STSCIC::SMM "On Duty SMM, EXT 5074" 16-OCT-1993 02:07:05.39 To: SMM CC: Subj: telnet.log HUBBLE SPACE TELESCOPE DAILY REPORT #994 PERIOD COVERED: 0800 EST Daily Status Report as of 288/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make two standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were executed as scheduled. One image was underexposed, and the other was overexposed. 1.2 Completed PC 5072 (WF/PC Cycle 2 Calibration: PC Flats Wide. 1) Both the Wide Field Camera and the Planetary camera were used to make Earth Flat calibration exposures using the wide and medium band filters. Four observations were received and displayed in the OSS. Two of the observations were reported to be okay, one was saturated, and one was underexposed. 1.3 Completed FGS 5057 (Extragalactic Astrometry and Astrophysics FT0-11, Part 1 fo 2) Fine Guidance Sensor #3 was used to make astrometric observations of of several different extragalactic objects. These observations will be used to determine the rotation of the reference frame used by the HIPPARCOS astrometry satellite. This proposal is tied to proposal 1013. The observation occurred during LOS and could not be monitored in real-time. No anomalies were reported. 1.4 Completed Three Sets of WFC 4848 (GTO Low Latitude Parallel Program - Cycle 3 High) The Wide Field Camera was used to make five high latitude medium-deep sky survey observations taken in parallel with the prime science observations. All five observations were received and displayed. They were reported to look okay. 1.5 Completed FGS 5059 (Seven Points of Light - Cycle 2 Calibration - Fixed) The Fine Guidance Sensors were used to make observations of standard star UPGREEN-69. The objective was to calibrate the single star interferometeric fringe pattern at seven different points in the central region of the astrometric FGS. The engineering link was not available and the observations were not monitored by the OSS. No anomalies were reported. 1.6 Completed FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) Fine Guidance Sensor #3 was used to make astrometric observations of the binary star system HD73900 to obtain its orbital elements. These observations are important since most of the estimates of stellar mass are made by observing elements of these systems. The astrometry was monitored and no problems were seen in the telemetry. 1.7 Completed GHRS/2 4541 (Cepheid Masses - Cycle 3 High) The High Resolution Spectrograph was used to make spectral observations of star HD172167 (aka, Vega). This star is being used as a wavelenght standard star for subsequent observations. The observations were executed as scheduled, and no problems were reported. 1.8 Completed WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) The Wide Field Camera was used to make an internal flat calibrations image. The image was reported to look normal. 1.9 Uplinked Updated Ephemeris Tables, per MOSES 2871 1.10 Completed FOS/BL 4058 (Mass Exchange Binaries (FOS 34) Augmentation: Cycle 2 Observation) The Faint Object Spectrograph (Blue Side) was used to make ultraviolet Rapid Mode spectropolarmetric observations of the eclipsing binary star system DQ-Hec. The white dwarf in the system has a rotational period of 71 seconds. The observations were monitored in the OSS, and the data quality was assessed. The target was located and the first observation was reported to be weak. The next observation had an overlight sum of 7000, and the data still remains on the science tape recorder. 1.11 Completed FOS/RD 5225 (Quasar Absorption Line Survey - Cycle 3 High - Part 3) The Faint Object Spectrograph (Red Side) was used to make two observations of quasar PKS0743_67 to learn more about the origin and evolution of gas clouds along the line of sight of this target. This is a continuation of proposal 2424. It is hoped to obtain a homogeneous sample of the QSO's spectra. Both observations were received and displayed. They had only noise, as was expected for this observations. No other anomalies were noted. 1.12 Completed WFC 4381 (HST Medium - Deep Survey - Cycle 3) The Wide Field Camera was used to make two set of medium-deep high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 10 Successful : 9 At 287/2156Z both FHST 1 and 2 failed to execute a full axis maneuver update. Subsequent acquisitions were successful. 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.2 Complete FOS/BL 4260 ()Extreme Stellar Winds and Post Main Sequence Evolution) 3.3 Complete Three Sets of WFC 4381 (HST Medium Deep Survey - Cycle 3) 3.4 Complete GHRS/2 4674 (Molecular Emissions in HH 47 - Cycle 3 High) 3.5 Complete WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.6 Complete WFC 4348 (Long Period Variables of M-81 - Cycle 3 Medium) 3.7 Complete Two Sets of FGS 4938 (Parallax of Hyades Cluster Members - Cycle 3 High) 3.8 Complete Two Sets of WFC 4848 (GTO Low Latitude Parallel Program Cycle 3 High) 3.9 Complete FGS 2565 (Linking HIPPARCOS to the Extragalactic Reference Frame Part 1 of 6) 3.10 Complete Two Sets of FOS/BL 4059 (Absolute Photometry Monitor) 3.11 Complete WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) 3.12 Complete PC 5073 (WF/PC Cycle 3 Calibration: PC Flats Wide. 2) 3.13 Complete PC 4712 (WF/PC Cycle 2 Calibration: Measles Monitor) 3.14 Complete FOS/RD/BL 4694 (Stellar Content of Galaxies and Globular Clusters) 3.15 Complete Four Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.16 Complete HSP/UV2 3798 (A search for Radial Pulsations in White Dwarfs) 3.17 Complete FOC/96 4846 (High Resolution Observations of Cataclysmic Variables - Cycle 3 High) 3.18 Complete WFC 4849 (GTO High Latitude Parallel Program - Cycle 3 High) 3.19 Complete PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Internal Simulation Type 10 (EVA#4) is scheduled for Tuesday Oct. 19, 1993 and will last from 8AM until 8PM. /DMH -END-